# Evidence for a Sharp CO Snowline Transition in a Protoplanetary Disk and Implications for Millimeter-wave Observations of CO Isotopologues

### v1

@@ 3 Application to HD 163296
The Herbig Ae star HD 163296 is surrounded by a disk that has been a favorite target of detailed study with the improving capabilities of millimeter interferometry for more than 25 years (e.g., Mannings & Sargent, 1997; Isella et al., 2007; Qi et al., 2011; Rosenfeld et al., 2013; Öberg et al., 2021). The N2H+ emission from this disk shows a characteristic bright, narrow ring with extended tenuous emission, which is well matched by the presence of a thick VIRaM layer (Qi et al., 2015) as found using the D’Alessio Irradiated Accretion Disk (DIAD) model (D’Alessio et al., 2006) that provided a consistent fit to the spectral energy distribution and overall shape of resolved millimeter dust continuum emission. This model features a refined vertical structure where dust settling is regulated by the transition location between small and large grains, $z=z_{big}$ in units of the gas scale height $H$. This grain regulation enhances the temperature disparity between the disk surface and interior. Larger dust grains, with lower infrared opacities, are concentrated near the midplane, resulting in less efficient heating compared to smaller particles. Consequently, the HD 163296 disk has a cold midplane with large grains at a significant scale height ($z_{big}=2H$). Near the midplane, the temperature remains approximately constant with height, effectively creating a thick VIRaM layer. Given the known CO snowline location around $\farcs$8 (Qi et al., 2015), this disk serves as an ideal example to study the sharp drop in CO column density across the CO snowline for a disk with a thick VIRaM layer.
@@ 3.1 ALMA Observations of CO Isotopologues Toward HD 163296
Figure 4: Radial profiles and derivatives of $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emission observed by ALMA toward HD 163296. The grey dashed line in each panel indicates the location of the CO snowline. The local minima in the derivatives for the optically thin tracers align closely with the CO snowline location.
@@ 3.2 $\mathrm{C^{18}O}$ 1-0 in Other MAPS Disks
For the MWC 480 disk, an intriguing feature is observed—a dip in the derivative of the C18O 1-0 radial profile at a radius smaller than $\farcs$3. However, this dip cannot be associated with the CO snowline, as thermal-chemical modeling of the MWC 480 disk places the CO snowline beyond 125 au ($\farcs$8) (Zhang et al., 2021a). This discrepancy suggests that other physical processes, such as chemical evolution, substructures, or line opacity effects, similar to those that affect the C18O 2-1 line for HD 163296, are influencing the emission profile in this disk.
Figure 5: Radial profiles and derivatives of $\mathrm{C^{18}O}$ 1-0 emission observed by ALMA toward MAPS sources GM Aur, AS 209, IM Lup, and MWC 480. The grey dashed line in the panel of GM Aur indicates the location of the CO snowline determined in Qi et al. (2019).
@@ 3.3 CO Depletion Factor
While Qi et al. (2015) fit the $C^{18}O$ 2-1 data visibilities by varying both the CO freeze-out temperature (and, consequently, the location of the CO snowline) and the CO depletion factor ($df$), our study focuses on evaluating the local emission profile change across the CO snowline for the optically thinner CO isotopologues with various $df$, avoiding the need to account for potential CO photodesorption by non-thermal processes beyond 200 au (Salinas et al., 2017). Since the CO snowline location has already been identified through the local minima in the derivatives of the emission profiles, as shown in Section 3, we explored models with depletion factors only, spanning a coarse logarithmic grid from 0.5 to 3.0 in steps of 0.5. To evaluate the relative flux change across the CO snowline, we normalized the flux profiles at a reference point near the snowline transition, specifically at $\farcs$65 (slightly interior to the snowline at $\farcs$8), to ensure sensitivity to the sharpness of the transition. Figure 6 shows the normalized flux profiles of $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emissions for models with a depletion factor of $df=5$ (as derived by Qi et al. (2015)) and $df=10^{2.5}$ (i.e., 320), compared with the observations. The model with a depletion factor of 320 matches the data more closely than the model with a depletion factor of 5, particularly for the optically thinner tracers $\mathrm{C^{17}O}$ 1-0, 2-1, and $\mathrm{C^{18}O}$ 1-0. The similarity in the shapes of the $\mathrm{C^{18}O}$ 2-1 flux profiles in both models and the observations indicates that this emission line is optically thick and largely insensitive to the depletion factor.
Figure 6: Normalized flux profiles of HD 163296 $\mathrm{C^{17}O}$ 1-0, 2-1 and $\mathrm{C^{18}O}$ 1-0, 2-1 emissions for the ALMA observations (blue) and for disk models with CO depletion factors of 320 (orange) and 5 (green). The model with a depletion factor of 320 matches the data better. The $\mathrm{C^{18}O}$ 2-1 emissions is not sensitive to the CO depletion factor.
@@ 4 Discussion
We have shown that emission from rare CO isotopologues provides a useful probe for the presence of a sharp CO snowline transition in disks with a thick vertical isothermal layer. We provided direct evidence for this feature in the disk around the Herbig Ae star HD 163296 from low J CO isotopologue emission resolved by ALMA observations. We next discuss implications of this thermal structure for CO distribution models of HD 163296, the general use of CO isotopologue emission for estimating total disk gas masses, and the feasibility of future observations in more disks.
@@ 4.1 Comparison of HD 163296 CO Distribution Models
The sensitivity of the low-J CO isotopologue lines to the CO snowline structure is underscored by comparing the emission resulting from different disk models. Figure 7 (left panel) shows the radial profile of the CO column density for three models of the disk around HD 163296: (1) the model with a depletion factor ($df$) of 5, as derived by Qi et al. (2015); (2) the model with a higher $df$ of 320, which provides a better match to the data, as demonstrated in Section 3.3; and (3) the model from the ALMA MAPS program, where the analysis of CO isotopologues suggested an enhancement in CO abundance interior to the snowline (Zhang et al., 2021a). The key region of interest in this comparison lies between 70 and 90 au, where the CO column density shows significant variations across the CO snowline. The model with $df=320$ demonstrates a much steeper decline in CO column density compared to either the $df=5$ model or the MAPS model. Specifically, the CO column density in the $df=320$ model decreases by more than a factor of 20 between 70 and 90 au, while the $df=5$ model shows a decrease by a factor of 5, and the MAPS model shows a decrease by only a factor of 2. This pronounced drop in the $df=320$ model provides clear evidence of a sharp CO snowline transition in the disk of HD 163296.
Figure 7: Left: Comparison of CO column density radial profiles for three different models of the HD 163296 disk: the MAPS model from Zhang et al. (2021a) (blue), the model with a CO depletion factor of 5 from Qi et al. (2015) (orange), and the model presented in this paper with a CO depletion factor of 320 (green). Reference CO column density values corresponding to $\tau=1$ for different CO isotopologue transitions based on RADEX calculations (see text) are marked in gray: $2\times 10^{20}\,\text{cm}^{-2}$ for C17O 1-0, $7\times 10^{19}\,\text{cm}^{-2}$ for C17O 2-1, $5.5\times 10^{19}\,\text{cm}^{-2}$ for C18O 1-0, and $1.9\times 10^{19}\,\text{cm}^{-2}$ for C18O 2-1. assuming ISM ratios of CO/C17O = 2005 and CO/C18O = 557. Right: Radial profiles of the normalized integrated flux of C17O 1-0 (blue) and C18O 1-0 (orange) emission observed with ALMA toward the disk of HD 163296. The profiles are normalized to their respective values at a radius of 70 au to facilitate comparison of the relative flux changes. Error bars indicate the observational uncertainties. The relative flux of C18O 1-0 remains consistently higher than that of C17O 1-0, suggesting that C18O 1-0 is optically thicker than C17O 1-0 across the CO snowline.
@@
We employed the DIAD model adopted from Qi et al. (2011, 2015), which directly addresses the degeneracy in vertical temperature structure that arises when models are constrained by the dust emission (e.g., the SED). The DIAD model accounts for the concentration of large grains near the midplane, enhancing the temperature gradient between the surface and interior layers. This configuration can result in a vertically isothermal region (VIRaM layer) near the midplane of some disks, influencing the CO distribution and its emission properties. In contrast, the Zhang et al. (2021a) analysis uses thermal-chemical models, such as RAC2D, to derive the dust temperature structure. While these models provide detailed chemical insights, they do not explicitly capture the vertically isothermal structure near the midplane. The difference in modeling approaches contributes to the different interpretations of the CO snowline transition in the HD 163296 disk across studies.
@@ 4.2 Effect on Disk Mass Estimates
As the models in Section 2 demonstrate, the thickness of the Vertically Isothermal Region at the Midplane (VIRaM) layer plays a critical role in shaping the vertical temperature distribution, which directly impacts the CO column density’s radial variation across the CO snowline. This, in turn, influences the emission of CO isotopologues and affects the derived disk mass.
Figure 8: Comparison of two disk models with different VIRaM layer thicknesses and disk masses. The left panel shows the CO column density profiles for Model 1 (thin VIRaM layer, 0.01 M, blue) and Model 2 (thick VIRaM layer, 0.04 M, orange). Model 2 exhibits a sharp drop in CO column density across the CO snowline, while Model 1 shows a more gradual decline. The right panel shows C17O 1-0 emission spectra for both models. The integrated flux of Model 1 is consistently higher than Model 2, despite Model 2 having higher CO column density inside the snowline, as emission from beyond the CO snowline, where the gas is optically thin, dominates at this (low) angular resolution.
@@ 4.3 Prospects with ngVLA
An important implication of our modeling is that high resolution observations of optically thin CO isotopologue line emission are needed to locate CO snowlines in disks with different thermal structures and to confront associated predictions. The combination of resolution and sensitivity will become even more important for disks around pre-main-sequence stars of lower luminosity than HD 163296 – the T Tauri stars – where the CO snowline is located closer to the star, typically $\sim 20-30$ au, and CO freeze-out is more significant. While it is clear that ALMA can effectively probe optically thin CO isotopologue lines in a large, warm, massive disk like the one around the Herbig Ae star HD 163296, even the relatively deep ($2-3$ hour) ALMA MAPS observations of disks around T Tauri stars resulted in tentative or modest detections of C17O 1-0 emission at $\sim 0\farcs 25$ resolution (Öberg et al., 2021). Substantially better observational capabilities than currently available with ALMA will be needed to obtain sufficient signal-to-noise ratio to extend the study of CO snowlines to a broader population of disks.
For sharp CO snowlines, detecting radial C/O variations becomes easier, offering a clearer understanding of the abundance and location of molecules in disks and their role in the formation of planetesimals and icy planets. The ngVLA is envisioned to improve on ALMA’s sensitivity by up to an order of magnitude for observations in the frequency range $35-116$ GHz where ngVLA and ALMA overlap. To date, most interest in ngVLA for protoplanetary disks has focused on milliarcsecond resolution observations of thermal dust emission to image terrestrial planet formation zones of disks, to reveal substructures associated with young planets (e.g., Ricci et al., 2018). However, the upper end of the ngVLA frequencies includes the ground state transitions of the suite of rare CO isotopologues that can provide optically thin tracers for a wide range of CO column densities at angular resolutions appropriate for CO snowlines in the large population of T Tauri stars located at the $\sim 140$ pc distance of nearby star-forming regions. Deep, resolved ngVLA observations to exploit these diagnostics. For example, the ALMA MAPS detection of C17O 1-0 in GM Aur at the $3.5\sigma$ level suggests the ngVLA could provide a resolved radial profile at sufficient signal-to-noise ratio to reveal the CO snowline location with a comparable investment of observing time. Observations of a large sample has the potential to provide a much more nuanced understanding of the factors that contribute to CO isotopologue emission from disks and the effects on inferred physical properties such as disk gas masses.
@@ 5 Conclusions
We use simple models of protoplanetary disks to demonstrate the importance of accounting for the disk thermal structure on the CO column density distribution, in particular the vertical extent of an isothermal region at the midplane, and we investigate the implications for observations of low J rotational lines of CO isotopologues at millimeter wavelengths. The main findings are:
@@
The development of the ngVLA, which will provide a substantial sensitivity improvement over ALMA for high resolution observations of the J=1–0 lines of CO isotopologues, will enable extending this approach to CO snowline studies to many more systems, in particular for a sample of the large population of T Tauri disks in nearby star-forming regions.
We thank the anonymous referee for a constructive report which improved the paper. C.Q. thanks K. Öberg, S. Andrews, and C. Espaillat for helpful discussions about the results presented in this paper. C.Q. acknowledges support for this work from the NRAO ngVLA Community Study program. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2016.1.00884.S and ADS/JAO.ALMA#2018.1.01055.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.

### v2

@@ 3 Application to HD 163296
The Herbig Ae star HD 163296 is surrounded by a disk that has been a favorite target of detailed study with the improving capabilities of millimeter interferometry for more than 25 years (e.g., Mannings & Sargent, 1997; Isella et al., 2007; Qi et al., 2011; Rosenfeld et al., 2013; Öberg et al., 2021). The N2H+ emission from this disk shows a characteristic bright, narrow ring with extended tenuous emission, which is well matched by the presence of a thick VIRaM layer (Qi et al., 2015) as found using the D’Alessio Irradiated Accretion Disk (DIAD) model (D’Alessio et al., 2006) that provided a consistent fit to the spectral energy distribution and overall shape of resolved millimeter dust continuum emission. This model features a refined vertical structure where dust settling is regulated by the transition location between small and large grains, $z=z_{big}$ in units of the gas scale height $H$. This grain regulation enhances the temperature disparity between the disk surface and interior. Larger dust grains, with lower infrared opacities, are concentrated near the midplane, resulting in less efficient heating compared to smaller particles. Consequently, the HD 163296 disk has a cold midplane with large grains at a significant scale height ($z_{big}=2H$). Near the midplane, the temperature remains approximately constant with height, effectively creating a thick VIRaM layer. Given the known CO snowline location around $\farcs$8 (Qi et al., 2015), this disk serves as an ideal example to study the sharp drop in the CO column density across the CO snowline for a disk with a thick VIRaM layer.
@@ 3.1 ALMA Observations of CO Isotopologues Toward HD 163296
Figure 4: Radial profiles and derivatives of $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emission observed by ALMA toward HD 163296. The gray dashed line in each panel indicates the location of the CO snowline. The local minima in the derivatives for the optically thin tracers align closely with the CO snowline location.
@@ 3.2 $\mathrm{C^{18}O}$ 1-0 in Other MAPS Disks
For the MWC 480 disk, an intriguing feature is observed: a dip in the derivative of the C18O 1-0 radial profile at a radius smaller than $\farcs$3. However, this dip cannot be associated with the CO snowline, as thermal-chemical modeling of the MWC 480 disk places the CO snowline beyond 125 au ($\farcs$8) (Zhang et al., 2021a). This discrepancy suggests that other physical processes, such as chemical evolution, substructures, or line opacity effects, similar to those that affect the C18O 2-1 line for HD 163296, are influencing the emission profile in this disk.
Figure 5: Radial profiles and derivatives of $\mathrm{C^{18}O}$ 1-0 emission observed by ALMA toward MAPS sources GM Aur, AS 209, IM Lup, and MWC 480. The gray dashed line in the panel of GM Aur indicates the location of the CO snowline determined in Qi et al. (2019).
@@ 3.3 CO Depletion Factor
While Qi et al. (2015) fit the $C^{18}O$ 2-1 data visibilities by varying both the CO freeze-out temperature (and, consequently, the location of the CO snowline) and the CO depletion factor ($df$), our study focuses on evaluating the local emission profile change across the CO snowline for the optically thinner CO isotopologues with various $df$, avoiding the need to account for potential CO photodesorption by nonthermal processes beyond 200 au (Salinas et al., 2017). Since the CO snowline location has already been identified through the local minima in the derivatives of the emission profiles, as shown in Section 3, we explored models with depletion factors only, spanning a coarse logarithmic grid from 0.5 to 3.0 in steps of 0.5. To evaluate the relative flux change across the CO snowline, we normalized the flux profiles at a reference point near the snowline transition, specifically at $\farcs$65 (slightly interior to the snowline at $\farcs$8), to ensure sensitivity to the sharpness of the transition. Figure 6 shows the normalized flux profiles of $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emissions for models with a depletion factor of $df=5$ (as derived by Qi et al., 2015) and $df=10^{2.5}$ (i.e., 320), compared with the observations. The model with a depletion factor of 320 matches the data more closely than the model with a depletion factor of 5, particularly for the optically thinner tracers $\mathrm{C^{17}O}$ 1-0, 2-1, and $\mathrm{C^{18}O}$ 1-0. The similarity in the shapes of the $\mathrm{C^{18}O}$ 2-1 flux profiles in both models and the observations indicates that this emission line is optically thick and largely insensitive to the depletion factor.
Figure 6: Normalized flux profiles of HD 163296 $\mathrm{C^{17}O}$ 1-0, 2-1 and $\mathrm{C^{18}O}$ 1-0, 2-1 emissions for the ALMA observations (blue) and for disk models with CO depletion factors of 320 (orange) and 5 (green). The model with a depletion factor of 320 matches the data better. The $\mathrm{C^{18}O}$ 2-1 emission is not sensitive to the CO depletion factor.
@@ 4 Discussion
We have shown that emission from rare CO isotopologues provides a useful probe for the presence of a sharp CO snowline transition in disks with a thick vertical isothermal layer. We provided direct evidence for this feature in the disk around the Herbig Ae star HD 163296 from low-$J$ CO isotopologue emission resolved by ALMA observations. We next discuss implications of this thermal structure for CO distribution models of HD 163296, the general use of CO isotopologue emission for estimating total disk gas masses, and the feasibility of future observations in more disks.
@@ 4.1 Comparison of HD 163296 CO Distribution Models
The sensitivity of the low-$J$ CO isotopologue lines to the CO snowline structure is underscored by comparing the emission resulting from different disk models. Figure 7 (left panel) shows the radial profile of the CO column density for three models of the disk around HD 163296: (1) the model with a depletion factor ($df$) of 5, as derived by Qi et al. (2015); (2) the model with a higher $df$ of 320, which provides a better match to the data, as demonstrated in Section 3.3; and (3) the model from the ALMA MAPS program, where the analysis of CO isotopologues suggested an enhancement in CO abundance interior to the snowline (Zhang et al., 2021a). The key region of interest in this comparison lies between 70 and 90 au, where the CO column density shows significant variations across the CO snowline. The model with $df=320$ demonstrates a much steeper decline in the CO column density compared to either the $df=5$ model or the MAPS model. In particular, the CO column density in the $df=320$ model decreases by more than a factor of 20 between 70 and 90 au, while the $df=5$ model shows a decrease by a factor of 5, and the MAPS model shows a decrease by only a factor of 2. This pronounced drop in the $df=320$ model provides clear evidence of a sharp CO snowline transition in the disk of HD 163296.
Figure 7: Left: comparison of CO column density radial profiles for three different models of the HD 163296 disk: MAPS model from Zhang et al. (2021a) (blue), model with a CO depletion factor of 5 from Qi et al. (2015) (orange), and model presented in this paper with a CO depletion factor of 320 (green). Reference CO column density values corresponding to $\tau=1$ for different CO isotopologue transitions based on RADEX calculations (see text) are marked in gray: $2\times 10^{20}\,\text{cm}^{-2}$ for C17O 1-0, $7\times 10^{19}\,\text{cm}^{-2}$ for C17O 2-1, $5.5\times 10^{19}\,\text{cm}^{-2}$ for C18O 1-0, and $1.9\times 10^{19}\,\text{cm}^{-2}$ for C18O 2-1, assuming ISM ratios of CO/C17O = 2005 and CO/C18O = 557. Right: radial profiles of the normalized integrated flux of C17O 1-0 (blue) and C18O 1-0 (orange) emission observed with ALMA toward the disk of HD 163296. The profiles are normalized to their respective values at a radius of 70 au to facilitate comparison of the relative flux changes. Error bars indicate the observational uncertainties. The relative flux of C18O 1-0 remains consistently higher than that of C17O 1-0, suggesting that C18O 1-0 is optically thicker than C17O 1-0 across the CO snowline.
@@
We employed the DIAD model adopted from Qi et al. (2011, 2015), which directly addresses the degeneracy in vertical temperature structure that arises when models are constrained by dust emission (e.g., the SED). The DIAD model accounts for the concentration of large grains near the midplane, enhancing the temperature gradient between the surface and interior layers. This configuration can result in a vertically isothermal region (VIRaM layer) near the midplane of some disks, influencing the CO distribution and its emission properties. In contrast, the Zhang et al. (2021a) analysis uses thermal-chemical models, such as RAC2D, to derive the dust temperature structure. While these models provide detailed chemical insights, they do not explicitly capture the vertically isothermal structure near the midplane. The difference in modeling approaches contributes to the different interpretations of the CO snowline transition in the HD 163296 disk across studies.
@@ 4.2 Effect on Disk Mass Estimates
As the models in Section 2 demonstrate, the thickness of the VIRaM layer plays a critical role in shaping the vertical temperature distribution, which directly impacts the CO column density’s radial variation across the CO snowline. This, in turn, influences the emission of CO isotopologues and affects the derived disk mass.
Figure 8: Comparison of two disk models with different VIRaM layer thicknesses and disk masses. The left panel shows the CO column density profiles for Model 1 (thin VIRaM layer, 0.01 M, blue) and Model 2 (thick VIRaM layer, 0.04 M, orange). Model 2 exhibits a sharp drop in the CO column density across the CO snowline, while Model 1 shows a more gradual decline. The right panel shows C17O 1-0 emission spectra for both models. The integrated flux of Model 1 is consistently higher than Model 2, despite Model 2 having higher CO column density inside the snowline, as emission from beyond the CO snowline, where the gas is optically thin, dominates at this (low) angular resolution.
@@ 4.3 Prospects with ngVLA
An important implication of our modeling is that high-resolution observations of optically thin CO isotopologue line emission are needed to locate CO snowlines in disks with different thermal structures and to confront associated predictions. The combination of resolution and sensitivity will become even more important for disks around pre-main-sequence stars of lower luminosity than HD 163296 – the T Tauri stars – where the CO snowline is located closer to the star, typically $\sim 20-30$ au, and CO freeze-out is more significant. While it is clear that ALMA can effectively probe optically thin CO isotopologue lines in a large, warm, massive disk like the one around the Herbig Ae star HD 163296, even the relatively deep ($2-3$ hr) ALMA MAPS observations of disks around T Tauri stars resulted in tentative or modest detections of C17O 1-0 emission at $\sim 0\farcs 25$ resolution (Öberg et al., 2021). Substantially better observational capabilities than currently available with ALMA will be needed to obtain sufficient signal-to-noise ratio to extend the study of CO snowlines to a broader population of disks.
For sharp CO snowlines, detecting radial C/O variations becomes easier, offering a clearer understanding of the abundance and location of molecules in disks and their role in the formation of planetesimals and icy planets. The ngVLA is envisioned to improve ALMA’s sensitivity by up to an order of magnitude for observations in the frequency range $35-116$ GHz where ngVLA and ALMA overlap. To date, most interest in ngVLA for protoplanetary disks has focused on milliarcsecond-resolution observations of thermal dust emission to image terrestrial planet formation zones of disks, to reveal substructures associated with young planets (e.g., Ricci et al., 2018). However, the upper end of the ngVLA frequencies includes the ground-state transitions of the suite of rare CO isotopologues that can provide optically thin tracers for a wide range of CO column densities at angular resolutions appropriate for CO snowlines in the large population of T Tauri stars located at the $\sim 140$ pc distance of nearby star-forming regions. Deep, resolved ngVLA observations to exploit these diagnostics. For example, the ALMA MAPS detection of C17O 1-0 in GM Aur at the $3.5\sigma$ level suggests the ngVLA could provide a resolved radial profile at sufficient signal-to-noise ratio to reveal the CO snowline location with a comparable investment of observing time. Observations of a large sample have the potential to provide a much more nuanced understanding of the factors that contribute to CO isotopologue emission from disks and the effects on inferred physical properties such as disk gas masses.
@@ 5 Conclusions
We use simple models of protoplanetary disks to demonstrate the importance of accounting for the disk thermal structure on the CO column density distribution, in particular the vertical extent of an isothermal region at the midplane, and we investigate the implications for observations of low-$J$ rotational lines of CO isotopologues at millimeter wavelengths. The main findings are:
@@
The development of the ngVLA, which will provide a substantial sensitivity improvement over ALMA for high-resolution observations of the J=1–0 lines of CO isotopologues, will enable extending this approach to CO snowline studies to many more systems, in particular for a sample of the large population of T Tauri disks in nearby star-forming regions.
We thank the anonymous referee for a constructive report, which improved the paper. C.Q. thanks K. Öberg, S. Andrews, and C. Espaillat for helpful discussions about the results presented in this paper. C.Q. acknowledges support for this work from the NRAO ngVLA Community Study program. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2016.1.00884.S and ADS/JAO.ALMA#2018.1.01055.L. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under a cooperative agreement by Associated Universities, Inc.